### v6

# Interactive Multiscale Modeling to Bridge Atomic Properties and Electrochemical Performance in Li-CO2 Battery Design

Abstract
Li-CO2 batteries are promising energy storage systems due to their high theoretical energy density and CO2 fixation capability, relying on reversible Li2CO3/C formation during discharge/charge cycles. We present a multiscale modeling framework integrating Density Functional Theory (DFT), Ab-Initio Molecular Dynamics (AIMD), classical Molecular Dynamics (MD), and Finite Element Analysis (FEA) to investigate atomic and cell-level properties. The considered Li-CO2 battery consists of a lithium metal anode, an ionic liquid electrolyte, and a carbon cloth cathode with Sb0.67Bi1.33Te3 catalyst. DFT and AIMD determined the electrical conductivities of Sb0.67Bi1.33Te3 and Li2CO3 using the Kubo-Greenwood formalism and studied the CO2 reduction mechanism on the cathode catalyst. MD simulations calculated the CO2 diffusion coefficient, Li+ transference number, ionic conductivity, and Li+ solvation structure. The FEA model, parameterized with atomistic simulations data, reproduced the available experimental voltage-capacity profile at 1 mA/cm2 and revealed spatio-temporal variations in Li2CO3/C deposition, porosity, and CO2 concentration dependence on discharge rates in the cathode. Accordingly, Li2CO3 can form large and thin film deposits, leading to dispersed and local porosity changes at 0.1 mA/cm2 and 1 mA/cm2, respectively. The capacity decreases exponentially from 81,570 mAh/g at 0.1 mA/cm2 to 6,200 mAh/g at 1 mA/cm2, due to pore clogging from excessive discharge product deposition that limits CO2 transport to the cathode interior. Therefore, the performance of Li-CO2 batteries can be improved by enhancing CO2 transport, regulating Li2CO3 deposition, and optimizing cathode architecture.
1 Introduction
Lithium-carbon dioxide (Li-CO2) batteries, with their high theoretical energy density and CO2 fixation capability, have emerged as a potential energy storage solution [1, 2]. These batteries provide a theoretical specific energy density of 1876 Wh kg-1, which is substantially higher than that of current lithium-ion battery systems [3, 4]. This high energy density makes Li-CO2 batteries particularly attractive for applications in aviation, aerospace, and electric vehicles, where lighter and more compact energy storage solutions are crucial [5]. The potential for electric vehicles with much longer ranges or considerably lighter batteries could revolutionize the automotive market [6, 5]. However, Li-CO2 batteries are still in early development stages and face challenges including reaction reversibility issues, low energy efficiency, and the need for improved cathode designs [7]. Ongoing research and development efforts focus on overcoming these limitations, with recent advancements showing promising results [4, 8, 9, 10, 11, 12, 13, 14]. The Li-CO2 batteries operate on a unique electrochemical principle involving the reversible reaction between lithium and carbon dioxide [1, 2] contributing to CO2 capture and utilization efforts [11, 15]. During the discharge process, lithium oxidizes at the anode, releasing lithium ions and electrons . Besides, at the cathode in the presence of solid catalysts [16, 4], CO2 undergoes reduction (), combining with lithium ions and electrons to form primarily lithium carbonate (Li2CO3) and carbon (C) . This discharge reaction typically occurs at a voltage of approximately 2.8 V [17]. The charging process reverses these reactions, decomposing the discharge products to regenerate lithium and CO2. However, due to the stability of Li2CO3, this process requires higher voltages, often exceeding 3.8 V [18]. This increased voltage requirement underscores one of several key challenges facing Li-CO2 battery technology. In particular, the formation of solid Li2CO3 and carbon can lead to electrode clogging and increased internal resistance [9, 19]. Moreover, the high charging voltages needed to decompose Li2CO3 significantly reduce the overall energy efficiency of the system [20, 21]. To address these issues, effective cathode catalysts are critical for enhancing reaction kinetics and minimizing overpotentials in Li-CO2 batteries [8, 22]. Nevertheless, unwanted side reactions, particularly electrolyte decomposition, can occur at these high voltages [7]. Furthermore, the inherent inertness of CO2 compounds the challenge, as it makes electrochemical reduction difficult and adversely affects both discharge voltage and energy efficiency [23, 2, 15].
In this work, we consider a Li-CO2 battery cell comprising a lithium anode, a hybrid ionic liquid electrolyte, and a porous cathode. The electrolyte features a unique composition of 1-ethyl-3-methylimidazolium tetrafluoroborate (EMIM-BF4) and dimethyl sulfoxide (DMSO) in a 2:3 ratio, supplemented with zinc iodide (ZnI2) and 1 M lithium bis(trifluoromethanesulfonyl)imide (LiTFSI) salt. This optimized electrolyte formulation, experimentally validated for superior performance [4], enables efficient ion transport and stability. The cathode utilizes carbon cloth coated with a novel transition metal trichalcogenide alloy catalyst (Sb0.67Bi1.33Te3), which significantly enhances CO2 reduction/oxidation kinetics and cycling stability [4]. While recent advancements in Li-CO2 batteries have focused on developing novel cathode materials and electrolytes [10, 24, 25, 11, 14, 13, 26], most computational studies have been limited to atomistic-level investigations. To address this gap, we introduce the first comprehensive multiscale framework that integrates quantum-scale simulations, Molecular Dynamics, and continuum modeling of Li-CO2 batteries. This approach uniquely connects atomic-scale phenomena, such as catalyst surface reactions and ion transport, to macroscopic cell performance, enabling holistic optimization of Li-CO2 battery systems. This strategy is also needed for other battery types [27, 28, 29, 30]. Our methodology advances the field by addressing critical knowledge gaps in the spatiotemporal evolution of parameters and the interdependencies between electrochemical processes at different scales, and it can be extended to the study of other battery systems.
At the Li-CO2 battery cell level, Finite Element Analysis (FEA) can effectively simulate charge/discharge cycles, enabling macroscopic analysis of complex processes within the battery cell. These processes include ion transport through porous electrodes, reaction kinetics, and their impact on cell performance [31, 32, 33]. However, FEA requires accurate electrode and electrolyte properties as input parameters, which can be derived from atomistic simulations.
To bridge the gap between atomic and cell-level simulations and to calculate the necessary parameters for FEA, we employed a closed-loop workflow from the atomic to the cell level, as summarized in Figure 1. We employ quantum mechanics (QM) and Density Functional Theory (DFT) calculations to determine partial charges, equilibrated geometries, and Molecular Dynamics Force Field (FF) parameters for battery constituents using small systems (number of atoms < 1000). Ab Initio Molecular Dynamics (AIMD) calculations are used to determine the electrical conductivity of Li2CO3 and Sb0.67Bi1.33Te3, based on DFT-equilibrated geometries and employing the Kubo-Greenwood formula [34, 35, 36]. AIMD is also used to study the surface reactions of Sb0.67Bi1.33Te3 as a cathode catalyst material. Large-scale MD simulations are conducted to investigate the transport properties of the liquid electrolyte. Specifically, we calculate the diffusion coefficients of Li+ and CO2, as well as the Li+ transference number, by examining the mean square displacement (MSD) of the particles over time using the Einstein equation [37], and the ionic conductivity using the Green-Kubo formula [38]. The atomic-scale insights gained from DFT/AIMD and MD simulations are then integrated into our FEA model, enabling a comprehensive analysis of cell-level performance. Our FEA simulations investigate various aspects of battery cell behavior, including specific capacity, discharge behavior, and voltage-capacity profiles, considering factors such as CO2 transport properties, cathode porosity, and the deposition of Li2CO3 and C.
The Li-CO2 continuum model (CM) of the Li-CO2 battery that we developed, supported by accurately calculated parameters from atomistic-level simulations, enables us to make predictions about porosity, deposited species quantities, and CO2 concentration variations—parameters that are difficult to assess through experimental methods alone. Furthermore, it allows us to evaluate these properties as a function of time, providing dynamic insights into the battery’s behavior at the cell level. This hierarchical, cross-validated workflow not only bridges the longstanding gap between atomistic understanding and macroscopic battery behavior but also enables targeted material screening and iterative optimization of cell components. The resulting predictive capability accelerates the discovery and engineering of Li-CO2 batteries with enhanced energy density, cycle life, and operational safety. Moreover, our methodology provides a generalizable template for the rational design of other advanced battery chemistries, where multiscale interactions govern performance and reliability. Ultimately, this multiscale computational strategy offers a powerful tool to guide experimental efforts, reduce development time, and unlock the full potential of emerging energy storage technologies.
2 Methods and computational details
2.1 Density Functional Theory Calculations for Structure Optimization
First-principles Density Functional Theory (DFT) calculations were performed using the Vienna Ab-initio Simulation Package (VASP) [39, 40] to optimize the crystal structures of Sb0.67Bi1.33Te3 and Li2CO3. The Sb0.67Bi1.33Te3 structure was constructed based on Sb2Te3 from the Materials Project database [41], and multiple configurations were optimized to identify the most energetically favorable structure. The Li2CO3 structure was also obtained from the Materials Project database. For surface interaction studies, the Perdew-Burke-Ernzerhof (PBE) functional within the generalized gradient approximation (GGA) was used, and dispersion interactions were accounted for using the D3 van der Waals correction [42]. These calculations are crucial for accurately modeling CO2 adsorption and reduction processes on the catalyst surface. Further computational details are provided in the Supplementary Information.
2.2 Ab Initio Molecular Dynamics (AIMD) and Kubo-Greenwood method for electrical conductivity calculation
Using optimized structures of Sb0.67Bi1.33Te3 and Li2CO3 compounds from our DFT calculation, we carried out the Ab Initio Molecular Dynamics (AIMD) simulations at 300K up to 1ps with NVT ensemble and 1fs time step to select 10 different frames in the 1ps trajectory data. Subsequently, we generated wavefunctions ( where and are band and k-point indices, respectively) data for the selected frames using DFT calculations as explained by Knyazev et. al. [43, 34] to compute the electronic conductivity based on Kubo-Greenwood (KG) formula. During the calculation of wavefunction data from VASP, we included 100% more unoccupied bands than occupied bands for both Sb0.67Bi1.33Te3 and Li2CO3 compounds to account for excited state wavefunctions and their corresponding energies. The dynamic electrical conductivity is represented as a coefficient of the current density and electric field :
where is the frequency of the external electric field . The important component relates to the energy absorbed by the electrons and is computed by the Kubo-Greenwood formula as [34, 35, 36]
Here, the sum is over all points in the Brillouin zone, band indices and and three spatial dimensions . is the frequency-dependent electrical conductivity. is the elementary charge. is the reduced Planck constant. is the electron mass. is the angular frequency. is the volume of the system. and are the Fermi-Dirac distribution functions for states and . and are the wavefunctions of states and . and i are the energies of states and . is the Dirac delta function. denotes a weight function at a particular point. The wave function is generated from our AIMD simulations at room temperature as explained above. is the Fermi function. In practical calculations, -function is written in terms of the Gaussian function:
where E is broadening of the -function. This technical parameter plays an essential role in controlling the algorithm in the Kubo-Greenwood formula, particularly at low frequencies.
2.3 Atomistic Molecular Dynamics Simulations (MD)
Classical Molecular Dynamics (MD) simulations were performed using LAMMPS [44] to investigate the electrolyte system relevant to Li-CO2 batteries. The simulated electrolyte consisted of DMSO:(EMIM+, BF) at a 3:2 ratio, 1 M (Li+, TFSI-), and included 6000 DMSO molecules, 1857 (EMIM+, BF) pairs, 566 (Li+, TFSI-), 14 ZnI2, and 14 CO2 molecules, totaling 113,142 atoms. Initial configurations were generated and energy minimized, followed by agitation using the Langevin and Nosé-Hoover thermostats. The system was equilibrated to ambient conditions (300 K, 1 bar) in stages using the Berendsen barostat, and transport properties were evaluated under experimental-like conditions in the NVT ensemble at 300 K. Further details on force field parameterization, cutoff schemes, and equilibration protocols are provided in the Supplementary Information.
2.4 Finite element model
A mathematical framework has been developed for modeling the behavior of prismatic Li-CO2 batteries, drawing inspiration from previous studies on Li-O2 and Li-ion batteries [45, 46, 47]. The model integrates key principles such as mass and current conservation, species transport phenomena, and reaction kinetics within the cathode and separator regions to provide deeper insights into the complex mechanisms that occur within the cell during operation. The Li-CO2 cell under investigation consists of several components: a lithium metal sheet serving as the negative electrode, a separator, a porous carbon cathode coated with Sb0.67Bi1.33Te3, and a liquid electrolyte that permeates the porous structures, as illustrated schematically in Figure 2. The used electrolyte composition is described in section 2.3. To facilitate electron transfer, current collectors are located at the rear of both electrodes. This configuration enables a thorough examination of the electrochemical and transport processes that govern battery cell performance and limitations. The main objective of this mathematical framework is to provide a comprehensive understanding of the intricate interactions among various components and processes within the Li-CO2 battery cell system.
2.4.1 Electrode Reactions
Cathode Reaction
The porous carbon electrode provides a site for the electrochemical reduction of carbon dioxide. During operation, CO2 from the external feed side dissolves in the electrolyte, moves through the pores of the positive electrode, and reacts with lithium ions at the active sites. The reaction considered at the positive electrode is [11]:
where and represent the equilibrium potentials for discharge and charge reactions, respectively [17, 18]. The reaction products (Li2CO3 and C) have limited solubility in the DMSO-based liquid electrolyte (0 mol/L for C and 0.0371 mol/L for Li2CO3). When their concentrations exceed these solubility limits, the electrolyte solution becomes supersaturated [48], leading to the deposition of Li2CO3 and C. This is followed by the nucleation and growth of these products in the liquid phase within the porous electrode. These deposited products significantly impact the battery’s performance by covering the active surface area, obstructing pathways for reactive species (Li+ and CO2), and inhibiting further reactions within the cathode, thereby limiting the extent of cell discharge.
Anode Reaction
2.4.2 Particle transport at the macroscopic level
The Nernst-Planck equation, a continuity equation for the time-dependent concentration of a chemical species, is used to describe the transport of chemical species within the battery cell:
In this context, represents the porosity of the electrode or the electrolyte region in the solution phase, denotes the flux, and indicates the volumetric production rate of species transitioning from the electrode material to the electrolyte within the porous structure. The total flux is composed of three components: diffusion, advection, and electromigration. This implies that the concentration is affected by an ionic concentration gradient , flow velocity v, and an electric field E.
In this equation, denotes the effective diffusion coefficient of the chemical species, represents the valence of the ionic species, represents the elementary charge, represents the Boltzmann constant, and represents the absolute temperature. The electric field may be further decomposed as:
where is the electric potential and is the magnetic vector potential. Therefore, the Nernst–Planck equation is given by:
The primary transport mechanisms in batteries, ionic migration due to electric fields and diffusion due to concentration gradients, are accurately described by the Nernst-Planck equation, which excludes time-varying magnetic effects [49]:
The moving boundaries theory of transference number () gives [50]:
where is the volume occupied by diffusing cations by the boundary in time , the Li+ concentration, F the Faraday constant, and the electric current.
By introducing the electrode surface :
Also, we can write as:
with the number of particles and the Avogadro number.
Then:
with, J the molar flux of cations, the current density in the electrolyte, and the Faraday constant.
| Symbol | Value | Description | Reference |
| Parameters used in the battery assembly | |||
| Lsep | 50 m | Thickness of Separator | This work |
| Lpos | 2 mm | Thickness of positive electrode | This work |
| Ac | 1 cm2 | Cross-sectional area | This work |
| 0.73 | Initial porosity of positive electrode | This work | |
| m | 0.4 mg | mass of active cathode material | This work |
| 0.5 | Porosity of separator | ||
| 1- | Initial active material solid fraction of cathode | This work | |
| cLi,0 | 1000 mol/m3 | Initial concentration of Li+ in electrolyte | This work |
| Parameters calculated using DFT/AIMD simulation | |||
| 1.9 105 S/m | Conductivity of positive electrode | This work | |
| 1.8 10 S/m | electrical conductivity of Li2CO3 film | This work | |
| Parameters calculated using MD simulation | |||
| D | 3.17 10-9 m2/s | Diffusion coefficient of Li+ in electrolyte | This work |
| D | 1.2 10-9 m2/s | CO2 diffusion coefficient | This work |
| 6.77 mS/cm | Conductivity of Li+ in electrolyte | This work | |
| t+ | 0.91 | Transference number of Li+ in electrolyte | This work |
| General parameters | |||
| i0,Li | 0.965 A/m2 | Exchange current density for anode | [51, 46] |
| dlnf±/dlnc | -1.03 | Activity dependence | [51, 46] |
| 2.8 V | Equilibrium potential for discharge | [17] | |
| 3.8 V | Equilibrium potential for charge | [18] | |
| rpos,0 | 25 nm | Particle radius in the positive electrode | [46] |
| a0 | 3/rpos,0 | Initial active specific surface area | [46] |
| n | 4 | Number of transferred electrons | [17] |
| c | 0.0371 mol/L | Solubility limit of Li2CO3 in DMSO | |
| cC,max | 0 mol/L | Solubility limit of Carbon in DMSO | |
| 2140 kg/m3 | Density of Li2CO3 | ||
| M | 73.89 10-3 kg/mol | Molecular weight of Li2CO3 | |
| 2.11 g/cm3 | Density of carbon | ||
| T | 300 K | Temperature | |
| Assumed parameters | |||
| ka | 1.11 10-15 m/s | Anodic reaction rate coefficient | Assumed |
| kc | 3.14 10-15 m7/s/mol2 | Cathodic reaction rate coefficient | Assumed |
| c | 4.8 mol/m3 | Initial CO2 concentration in positive electrode | Assumed |
| Variable | Value | Boundary |
| Total current density | iapp | 3 |
| c | c | 3 |
| 0 V | 1 | |
| 0 | 2 | |
| 0 | 1, 3 |
| Variable | Value | Domain |
| 2.8 V | Porous cathode | |
| 0 V | Separator | |
| 0 V | Porous cathode, Separator | |
| cLi,0 | Porous cathode, Separator | |
| c | Porous cathode | |
| c | Porous cathode | |
| cC,max | Porous cathode | |
| Porous cathode |
The primary interest in battery modeling is electrochemical processes rather than fluid dynamics, so the advection term is considered less critical [52]. Furthermore, the porous structure limits large-scale fluid motion, further reducing the importance of advection. By neglecting the Advection term, the Nernst–Planck equation for Li+ and CO2 can be expressed as:
The effective diffusion coefficient of a species i (i=Li+, CO2) was calculated using the Bruggeman equation [32, 33]:
The form, in the Bruggeman relation, emerges from the mathematical treatment of a system of spherical particles dispersed in a continuous medium [53]. This form has been found to provide a reasonable empirical fit approximation for many real porous systems, especially at higher porosities [53, 54]. The diffusion coefficient (i=Li+, CO2) is calculated from our MD simulation.
2.4.3 Equations of battery electrochemical properties
The surface density of the liquid phase represents the function of the electrolyte concentration , time , the ionic conductivity , the absolute temperature , the activity coefficient , the universal gas constant , the cathode porosity and electrode solution-phase potential [45, 46].
The solid phase current density is the function of the solid-electrode potential and the electronic conductivity of solid-phase mixture [45, 46],
denotes the starting porosity of the cathode. It is important to mention that the effect of deposits (Li2CO3 and C) with low conductivity on was not taken into account, and it was assumed that would not change even as the battery reactions occurred. This assumption is primarily justified by the demonstrated structural and chemical stability of the catalyst under cycling conditions [4]. In our continuum model, we further decouple electron transport through the deposited Li2CO3 and carbon byproducts, treating their intrinsic conductivities separately from that of the catalyst. This approach ensures that the catalyst’s role in maintaining electronic percolation pathways is accurately represented. However, it is important to acknowledge the potential limitations of this simplification. If the catalyst were not stable-for example, if it were susceptible to pulverization or surface passivation-its conductivity could decrease during cycling. Such changes might result in localized current hotspots or capacity fade, which would not be captured by the current model. Therefore, while the assumption is reasonable for stable catalyst systems, it may not be universally applicable to all materials or operational conditions.
Conservation of charge
These equations demonstrate that the transfer current per unit electrode volume corresponds to the electrode chemical reaction rate, where is a symbol representing a species participating in the electrochemical reaction, and are the charge number and the stoichiometric coefficient of species , respectively, is the number of electrons transferred in the reaction, is the specific interfacial area of the pore per unit electrode volume, and is the average transfer current. In the present work, only one electrochemical reaction is considered: the formation of /carbon (/C) inside the porous cathode. In practical cells, battery side reactions can occur during cycling alongside the main electrode reactions. This is reflected by the superficial production rate of a species from the cathode’s solid phase to the pore solution, given by Faraday’s law as a function of the local transfer current density between the cathode/electrolyte interface.
Rate expressions
For modeling the cathode’s electrochemical reaction, a modified version of the Butler-Volmer equation is applied in the model using two rate coefficients. This modification is necessary because the cathode reaction presented in equation 4 depends on both the concentrations of Li+ and CO2, as well as the concentration of Li2CO3/C [55]:
where is local transfer current density between electrode and electrolyte interface, is the molar concentration of species at the cathode surface, and are the anodic and cathodic rate constant, respectively, is a symmetry factor equal to 0.5, is surface or activated overpotential for reaction at cathode, and are the voltage drop and the electrical resistivity across /C film formation, respectively, is the volume fraction of solid /C, and is the theoretical open-circuit potential for reaction.
The properties of the /C film, specifically its electrical resistivity , thickness (), and conductivity (), are related by the equations below:
where represents or C, while , , , , and refer to the film thickness of , the change in porosity of the cathode caused by , the molar mass of , the specific area, and the density of , respectively.
The electrochemical reaction at the lithium metal anode is given by the general Butler-Volmer equation as follows [45, 46]:
where is the exchange current density for the anode, is the surface or activated overpotential for the reaction at the anode, and the other parameters are as described above.
The specific area of the electrode/electrolyte interface in Equation 31 is decreased by the morphology and dynamic change in porosity due to /C solid passivation during the electrochemical reaction. This effective local surface area per unit volume of the electrode can be commonly written by a geometric relation [45, 46]:
where the initial specific surface area, and are the volume fraction of solid /C and initial electrode porosity, respectively.
Transport through /C layer
During discharge, the formed /C creates an additional barrier for Li+ and diffusion from the porous medium to the cathode surface.
This process follows Fick’s law [46], where the flux is proportional to the concentration gradient between the bulk electrolyte and the active surface, sustained by ongoing electrochemical reactions consuming reactants.
where is the effective diffusion coefficient of species across the /C film, and is the thickness of the film. represents the molar concentration of species in the electrolyte, while denotes the molar concentration of species specifically at the cathode surface. The effective diffusivity for both and CO2 reactants is used to describe how the pores are deposited by the /C formation; this can be described by the Bruggeman relation [32, 33]:
The volume fraction of the solid /C and the thickness of the film () can be determined from the cathode volume balance as:
In this context, and represent the molar mass and density of , respectively. The term denotes the volume fraction of the initial solid phase of the cathode material, while indicates the radius of the particles within the electrode.
Initial conditions/Boundary conditions
The governing equations consist of partial differential equations that include time and space as independent variables. The dependent variables were simulated using COMSOL Multiphysics 6 Finite Element Analysis software to examine their spatial and temporal changes. The battery cell was segmented into three regions: the anode, separator, and porous cathode. An extra fine mesh was selected for enhanced accuracy. To reduce computation time, a variable-step solver was employed with a relative tolerance of 1 10-4. This solver decreases the total number of time steps required for the calculation. The parameters used in the simulations are presented in Tables 1, 2, and 3. The applied current density, iapp, was considered positive when directed from the cathode to the external circuit (Figure 2). The total current density passing through boundary 3 was designated as iapp. The Li+ flux at boundaries 1 and 3 (JLi) and CO2 flux at boundary 2 (J) were set to zero. Additionally, the boundary condition ci = 0 was applied to the relevant boundaries (see Table 2). The parameter c was assumed to be equal to the saturation concentration, c, at boundary 3. The boundary conditions are outlined in Table 2. Discharge simulations were conducted with applied current densities (iapp) ranging from 0.1 to 1 mA/cm2, defined as positive in the direction from the cathode to the external circuit.
3 Results and discussion
3.1 Electrical conductivity from DFT, AIMD, and Kubo-Greenwood formula
The calculated dynamic electrical conductivities for Sb0.67Bi1.33Te3 and Li2CO3 are shown in Figure 3 (a1) and (b1) across a broad frequency range, and in Figures 3 (a2) and (b2), which focus on low frequencies to enable extrapolation to 0 Hz, with E values ranging from 0.22 eV to 0.46 eV. Static electrical conductivities at zero frequency were determined by extrapolating the low-frequency data in Figures 3 (a2) and (b2). The calculated conductivities are within the experimental order of magnitude. E values between 0.34 eV and 0.42 eV were selected because they show convergence with experimental conductivity data within this range. Furthermore, varying the electrical conductivity within the computational uncertainty does not significantly affect the predicted battery properties in the FEA. The static electrical conductivities calculated for different values of E show the conductive behavior of Sb0.67Bi1.33Te3 (a conductivity of the order of S/m) and insulative behavior of Li2CO3 (a conductivity of the order of S/m). Our computational analysis reveals a static electrical conductivity of S/m for Sb0.67Bi1.33Te3, indicating its strong electrical conducting properties. This finding aligns with the well-established characteristics of bismuth telluride (Bi2Te3) and its alloys, which are extensively employed in thermoelectric devices due to their good electrical conductivity [56, 57]. The high conductivity can be attributed to the semiconductor nature of Bi2Te3 and the antimony (Sb) doping effect, which augments the carrier concentration [58]. Experimental investigations on Bi2Te3 have reported electrical conductivity values as high as S/m at 323 K [58], further corroborating the computational results. Our calculations reveal that Li2CO3 exhibits a static electrical conductivity of S/m at 300 K, confirming its highly insulating nature. This result closely aligns with experimental measurements of S/m at 393 K [59]. The slight difference between calculated and experimental values is consistent with the expected increase in conductivity at higher temperatures, underscoring the reliability of our computational approach. The insulating nature of Li2CO3 can be attributed to its ionic nature and the strong ionic bonds that hinder electron mobility [60]. In Li2CO3, Li+ ions act as the primary charge carriers rather than electrons, a characteristic typical of ionic compounds. The low conductivity of Li2CO3 presents challenges to battery performance, including high internal resistance and passivation of the electrode surface [61, 62].
The calculated electrical conductivities are used in the continuum model to characterize electron percolation in the cathode and the deposited Li2CO3 during the battery charge/discharge cycle.
3.2 Surface reaction of Sb0.67Bi1.33Te3 as cathode catalyst
To comprehensively understand the formation of Li2CO3 and C on the Sb0.67Bi1.33Te3 catalyst surface, we calculated the Gibbs free energy () changes for various elementary reaction steps on this surface as shown below (steps 1-5) similar to the calculations done in Ref. [4]. Here, we used a slightly bigger supercell (as shown in Figure 3 (a1)) compared to the work done in Ref. [4]. The reaction pathway revealed a sequence of intermediate stages, beginning with the adsorption of CO on a Bi site with a Bi-O distance of 3.5 Å, which was marginally endergonic ( eV). The coupled addition of Li+ and an electron to the adsorbed CO formed the LiCO intermediate, a highly exergonic step ( eV). Subsequent interactions between two LiCO intermediates produced Li2CO and CO, with a of eV. Further, the conversion of Li2CO and CO into Li2CO3 and CO∗ ( eV) marked a pivotal step in the pathway. Finally, the generation of carbon was facilitated through the reaction of CO and CO2, an energetically uphill step ( eV) that also produced CO on the surface. These results, illustrated in Figure 4, highlight the essential role of coupled electron-cation transfer processes in the growth of discharge products. The results in this paper are consistent with the work reported in Ref. [4]. The Gibbs free energy of the reaction was calculated using the following equation:
where ZPE is the zero-point energy, is the heat capacity, is the entropy, is the temperature, and is the electronic energy.
In the calculation of Gibbs free energy profiles, full corrections including zero-point energy (ZPE) and entropy were applied explicitly to the initial gas-to-surface adsorption step of CO2, where such contributions are known to be significant due to the loss of translational and rotational entropy. For subsequent steps involving transformations between adsorbed intermediates (e.g., Li2CO3, Li2CO2, Li2CO3, and C–CO3 species), we assumed that ZPE and entropy effects largely cancel out, as these involve similar bonding environments and restricted surface-bound species. This approach is widely used in heterogeneous catalysis modeling and has been shown to introduce minimal quantitative error, while preserving the qualitative accuracy of energy landscapes and mechanistic conclusions [63]. Accordingly, our analysis emphasizes the dominant electronic energy contributions for these surface processes, which govern the reaction thermodynamics and barrier-lowering effects central to catalyst performance. The potential-dependent reaction energies were calculated based on the computational hydrogen electrode (CHE) model with reference to the Li/Li+ electrode [64, 65]. The potential-dependent reaction energies profile is shown in Figure 4, and the respective reaction energies of each step are shown in Table 4.
To level up the electrochemical step, a potential was applied to the formation of LiCO2. As can be seen in Figure 4, this potential occurs at 2.3 V versus Li/Li+ electrode [4].
The complete 5-step reaction pathway can be summarized as follows:
| (Step-1) | ||||
| (Step-2) | ||||
| (Step-3) | ||||
| (Step-4) | ||||
| (Step-5) |
The strong exergonic nature of steps 2-4 suggests that these processes occur readily, while the endergonic nature of the final carbon formation step aligns with the observed preferential adsorption of carbon on the Sb0.67Bi1.33Te3 catalyst surface. The Li2CO3 is formed in step 4, but this step also releases CO∗ and a free surface site (∗). This suggests that Li2CO3 formation occurs near, but not directly on, the catalyst surface, reducing the cathode porosity and possibly clogging the cathode pores.
These insights into the reaction mechanism have significant implications for catalyst design and overall battery performance. The strong affinity of carbon for the catalyst surface, coupled with the energetically unfavorable step of formation, suggests that catalyst degradation through carbon accumulation could be a concern over long-term battery operation. This highlights the need for strategies to mitigate carbon build-up, such as developing self-cleaning catalysts or implementing periodic regeneration protocols.
Furthermore, the spatial separation of Li2CO3 formation from the immediate catalyst surface underscores the importance of electrolyte engineering in Li-CO2 battery design. Since Li2CO3 formation appears to occur away from the catalyst, the electrolyte’s ability to facilitate Li2CO3 formation, dissolution, and transport becomes crucial.
The continuum model considers the formation of Li2CO3 and carbon in the cathode. We assume that the catalyst increases the reaction rate of Li2CO3/C formation and use a reaction rate constant on the order of 10-15 m7 s-1 mol-2. This value is higher than the typical value assumed for Li2O2 formation in Li-O2 batteries with non-catalyzed carbon cathodes, which is on the order of 10-17 m7 s-1 mol-2 [46].
| Reaction Step | Reaction Steps | (eV) |
| Step-1 | 0.002 | |
| Step-2 | -2.3 | |
| Step-3 | -1.2 | |
| Step-4 | -1.7 | |
| Step-5 | 2.8 |
3.3 Li-CO2 battery electrolyte transport properties and Li+ solvation structure from large scale MD simulation.
This section presents the results of MD simulations that elucidate the atomic-level dynamic and solvation properties of the Li-CO2 battery electrolyte depicted in Figure 5 (a). These simulations provide microscopic insights into the electrolyte’s behavior, bridging the gap between the atomic and macroscopic transport properties in our multiscale approach.
3.3.1 Li-CO2 battery electrolyte transport properties
The primary goal is to investigate the electrolyte’s dynamic properties governing ion and CO2 transport mechanisms, including ionic conductivity, cation transference number, and CO2 diffusion coefficient, as presented in Table 5. These properties directly impact the Li-CO2 battery’s electrochemical performance.
The cations’ transference number is assessed by [66]:
Where the self-diffusion coefficients are determined from the mean-square-displacement of the center of mass presented in Figure 5 (b), calculated from the MD simulations as follows [67]:
In the expression above, represents the temporal position of a random walker (ion or CO2 molecule). Here, indicates the time, and is the initial time at which the random walker begins to move. The diffusion coefficients are extrapolated from the normal diffusion regime characterized by:
where is the crossover time to the normal diffusion regime.
The Green-Kubo (GK) approach [38] that takes the autocorrelation with respect to the ionic current in the electrolyte is used to calculate the ionic conductivity :
with,
Here is the elementary charge, represents the charge number (valence) of ion , is the velocity of ion , is temperature, is the Boltzmann constant, is the volume of the simulation box, and is the number of ions.
| Parameter | Value | Unit | Symbol |
| Diffusion coefficient of Li+ | 3.17 | 10-9 m2s-1 | D |
| Ionic conductivity of the electrolyte | 6.77 (7.4)a | mS cm-1 | |
| Transference number of Li+ | 0.91 (0.88)a | — | |
| Diffusion coefficient of CO2 | 1.2 (1.16)b (1.14)c | 10-9 m2s-1 | D |
The estimated diffusion coefficient of CO2 in DMSO at 25 C is approximately 1.16 10-9 m2/s using the Stokes-Einstein-Gierer-Wirtz (SEGWE) relation and 1.14 10-9 m2/s using the Einstein equation [69]. These values are consistent with our MD calculation result of 1.2 10-9 m2/s presented in Table 5, further validating our findings.
The calculated diffusion coefficients, transference number, and ionic conductivity are directly incorporated into the FEA component of our multiscale modeling framework. These atomistically derived transport properties serve as critical input parameters for the FEA, enabling us to accurately simulate the intercalation and transport of CO2 and ions during battery charge/discharge cycling at the cell level.
3.3.2 Li+ solvation structure from large-scale MD simulation
The structural analysis was conducted utilizing the Radial Distribution Function (RDF), g(r), and Coordination Number (Nc). The RDF quantifies the spatial distribution of particles relative to a reference particle, providing insights into the local structural organization that affects the transport properties in the electrolyte.
The RDF is mathematically expressed as:
with the particle density of type at a distance around particles , and the particle density of type averaged over all spheres around particles with radius . Here, we chose =12Å.
The primary peak in g(r) corresponds to the nearest neighbors, indicating the most probable interparticle distance and interaction strength. This relationship is described by , where , and is the interaction potential. Subsequent peaks in the RDF elucidate higher-order neighbor relationships, offering insights into medium-range order, coordination numbers, and packing efficiency. In isotropic, homogeneous systems, g(r) asymptotically approaches unity at large distances, signifying a uniform particle distribution.
The coordination number, Nc, is defined as the number of neighboring particles within a specified cutoff distance of the pair interaction potential from a central particle. It is mathematically represented as:
where denotes the distance at which the interparticle interaction potential approaches zero, represents the bulk density, and is the average number density of particles at a given distance r, related to the RDF by = g(r).
The coordination number as a function of distance, N(r), is expressed as:
where r represents an arbitrary distance from a reference particle.
We present the average Radial Distribution Function () and Coordination Number () profiles for BF, TFSI-, and EMIM+ ions, as well as DMSO solvent, surrounding a single Li+ ion. Additionally, we analyze the and between molecules of the same type, represented by their center-of-mass atoms. Li+ ions preferentially coordinate with molecules containing highly electronegative atoms, i.e., atoms with a negative partial charge. The flat curves in Figure 5 (e) indicate a weak interaction between Li+ and EMIM+. Consequently, Li+ primarily coordinates with BF, TFSI-, and DMSO. According to Figure 5 (c), (d), and (f), Li+ ions interact strongly with BF, TFSI-, and DMSO. The most probable distances between Li+ and the constituent atoms are: BF (Li+-F 2.47 Å, Li+-B 3.22 Å), TFSI- (Li+-O 2.57 Å, Li+-F 2.85 Å), and DMSO (Li+-O 2.45 Å). The Li+ cations in the electrolyte are more solvated by BF and DMSO than TFSI-, as evidenced by the higher peaks in the and the corresponding coordination number curves of Li+-BF. BF and DMSO are closer to the Li+ ions than TFSI- ions, playing a crucial role in the dissociation of LiTFSI salt. However, TFSI- can also access the first coordination shell of Li+ due to strong interactions between its O and F atoms and Li+. The presence of both BF and TFSI- facilitates favorable coordination between Li+ and F, potentially leading to the formation of a Li-F solid electrolyte interphase (SEI) that protects the Li-metal anode. Interactions between molecules of the same type are weak due to electrostatic repulsion, resulting in well-dispersed molecules throughout the simulated electrolyte. The average first shell coordination number values for a tagged Li+ cation were determined using a threshold distance of 3 Å, based on the range of the first peaks. Key findings from the coordination number analysis include: Li+-Li+ ) and ), Li+-BF (, ), Li+-DMSO (), and Li+-TFSI- ( and ).
The high ionic conductivity (Table 5) of the simulated electrolyte is derived from several key factors revealed by the solvation structure and interaction analyses (Figure 5). Li+ cations are preferentially solvated by BF and DMSO, promoting LiTFSI salt dissociation and increasing free charge carriers. Weak Li+-TFSI- interactions facilitate easier Li+ transport. The presence of BF and TFSI- creates Li+ conducting channels through favorable Li+-F coordination. Well-dispersed ionic species due to weak same-type molecular interactions enhance efficient ion transport. Optimal coordination numbers (e.g., for BF, for DMSO) allow efficient Li+ transport while maintaining sufficient solvation. Low Li+-Li+ coordination () minimizes clustering, promoting ion mobility. Multiple solvating species create a dynamic environment, facilitating rapid solvating molecule exchange and enhancing Li+ mobility. The preferential solvation of Li+ by BF and DMSO allows for more free movement compared to the larger TFSI- anions, resulting in a higher proportion of the ionic current carried by Li+ and thus increasing the Li+ transference number.
3.4 Cell performance from Finite Element Analysis
The simulated voltage-capacity curves of the Li-CO2 battery cell during charge/discharge cycles are shown in Figure 6 (a). These curves were generated through Finite Element Analysis (FEA) using input parameters derived from atomistic simulations (Figure 3, Table 5) and additional material properties detailed in Tables 1, 2, and 3. The FEA results were validated against experimental data collected at a current density of 1 mAh/cm2, as described in the Supporting Information (SI). A strong correlation between simulated and experimental voltages is observed. During discharge, the voltage profile exhibits three distinct regions: (1) an initial sharp voltage drop due to kinetic overpotential () [70], (2) a linear decline attributed to resistive losses () across the growing Li2CO3/C film (average plateau potential: 2.65 V), (3) a "sudden death" characterized by rapid voltage collapse after 5000 mAh/g. The latter two phenomena, previously described for Li-O2 batteries as charge-transport limitations through Li2O2 films [71], exhibit analogous behavior in Li-CO2 systems. During charging, the cell maintains a stable potential of approximately 4.2 V until reaching 6000 mAh/g, followed by a steep voltage rise. The system achieves a specific capacity of 6200 mAh/g (0.48 mAh/cm2).
Figure 6 (b) illustrates the impact of applied current density of 1 mA/cm2, 0.5 mA/cm2, 0.25 mA/cm2, and 0.1 mA/cm2 on the simulated discharge curves of Li-CO2 battery cell and Figure S1 shows the same curves for current densities of 0.2 mA/cm2, 0.3 mA/cm2, 0.4 mA/cm2, 0.6 mA/cm2, 0.7 mA/cm2, 0.8 mA/cm2 and 0.9 mA/cm2. The discharge capacity exhibits a substantial reduction, decreasing from 81,570 mAh/g at a low current density of 0.1 mA/cm2 to 6,200 mAh/g at the high current density of 1 mA/cm2. This observation aligns with prior experimental results, indicating that current density has a significant influence on cell capacity [72]. Furthermore, the discharge voltage plateau decreases as the current density increases (Figure 6 (b)). The capacity decrease at elevated discharge rates can be attributed to limitations in CO2 transport through the cathode, which becomes saturated with electrolyte and unable to sustain the electrochemical reaction. Consequently, as the discharge rate increases, CO2 reduction is confined to a localized area near the cathode-current collector interface. Furthermore, the rapid reduction in porosity due to Li2CO3 and C deposition on the active area’s surface further impedes CO2 transport into the cell, preventing complete utilization of the electrode’s porosity.
Figure 6 (c) demonstrates that the capacity of the studied Li-CO2 battery cell decreases exponentially with increasing current density, which can be described by the following equation [73]:
where: = 7500 mAh/g is the baseline capacity, = 110000 mAh/g is a scaling factor, = 0.157 is a characteristic current density. At low current density (i 0.157 ), the capacity tends towards its maximum value of 117500 mAh/g. This is the sum of the baseline capacity () and the additional capacity represented by . As the current density increases, there is a rapid decrease in the additional capacity provided by , resulting in a decrease in the total capacity . The parameter determines the rate at which the capacity decreases with increasing current density. A smaller would mean a faster decrease in capacity for a given increase in current density, while a larger would indicate a slower decrease. As the current density increases further, the capacity approaches the baseline value of 7500 mAh/g. This suggests that at high current densities, the battery’s capacity is primarily limited to its baseline capacity, as the contribution from the exponential term becomes negligible. The exponential decrease in capacity with increasing current density highlights a trade-off between power output and energy capacity [74]. At higher current densities (i 0.157 ), the battery can deliver more power, but at the cost of reduced capacity. This is a critical consideration in applications where both high power and high capacity are required. The connection between cell voltage () and current density () during the discharge of the studied Li-CO2 battery cell in Figure 6 (d) is represented by the formula [73, 75]:
where V and V. The parameter represents the slope of the voltage change with the logarithm of the current density/(mA/cm2), while represents the voltage at a reference current density of 1 mA/cm2. This formula indicates that as the current density increases, the cell voltage decreases logarithmically, a common characteristic in electrochemical cells due to increased polarization and decreased voltage at higher current densities [73, 75].
Figure 7 summarizes the influence of key parameters on the voltage–capacity profiles of Li-CO2 batteries at a current density of 1 mA/cm2. To investigate the influence of key parameters on the voltage-capacity profiles (Figure 7), we performed a parameter sensitivity analysis using FEA to solve the model equations described in Section 2.4. The base case parameters, as provided in Tables 1, 2, and 3, were used as the reference. Each parameter was varied over a specified range while keeping the others at their base values. Specifically, the initial porosity () was varied from 0.6 to 0.9 in steps of 0.05, the cathodic reaction rate constant () from 1 10-19 to 1 10-14 m7/s/mol2 (logarithmically spaced), the initial CO2 concentration (c) from 2.8 to 6.8 mol/m3, the cathode thickness () from 0.05 to 0.3 mm, the ionic conductivity () from 1 to 10 mS/cm, and the CO2 diffusion coefficient () from 0.6 10-9 m2/s to 1.6 10-9 m2/s. For each set of parameters, the discharge process was simulated at a constant current density of 1 mA/cm2. As shown in Figure 7 (a), increasing the initial porosity () leads to higher battery capacity by expanding the available surface area and enhancing CO2 accessibility, thereby promoting more efficient electrochemical reactions. Figure 7 (b) highlights the critical role of the cathodic reaction rate constant (); the voltage–capacity profile exhibits high sensitivity to variations in , emphasizing the necessity of effective catalytic materials for optimal battery performance. In Figure 7 (c), it is evident that higher initial CO2 concentration () results in increased capacity and a more stable voltage plateau, attributed to improved reactant availability at the cathode. Figure 7 (d) demonstrates that increasing the cathode thickness () from 0.05 mm to 0.1 mm substantially enhances capacity due to the greater active volume; however, further increases beyond 0.1 mm do not provide additional capacity gains, likely due to diffusion limitations. Figure 7 (e) shows that electrolyte ionic conductivity () is a critical factor, with values above 2 mS/cm required to achieve high capacities (>4000 mAh/g) and improved voltage stability. Figure 7 (f) indicates that a higher CO2 diffusion coefficient () further enhances both capacity and voltage plateau by facilitating more efficient CO2 transport to active sites. Notably, the same trends are observed at lower current densities; Figure S2 depicts these effects at a current density of 0.4 mA/cm2. Taken together, these results underscore the importance of simultaneously optimizing electrode structure, electrolyte properties, and catalytic activity to maximize the performance of Li-CO2 batteries.
Figure 8 (a1) and (a2) illustrate how porosity () and volume fraction of Li2CO3/C deposited species () change over space and time with current densities of i=1 mA/cm2 and i=0.1 mA/cm2, respectively. Figure S3 presents a comparison of these parameters at i=0.1, i=0.25, i=0.5, and i=1 mA/cm2. For high current density (i=1 mA/cm2), porosity decreases gradually, primarily near the CO2 feed side of the cathode (X/L=1), from its initial value of to 0.48 at the end of discharge. This decrease is due to the formation of a thin film (flake-like) of Li2CO3/C [76], represented by the gradual increase of from 0 to 0.25 near the cathode feed side. This film formation leads to electrode clogging, impeding efficient CO2 diffusion into the cathode. For low current density (i=0.1 mA/cm2), Li2CO3/C forms larger structures [76], evidenced by the gradual increase of across a wider region of the cathode from 0 to 0.64. This results in a distributed porosity decrease from 0.73 to 0.09, initially allowing more CO2 transport into the cathode. However, as discharge progresses, the substantial porosity decrease hinders CO2 diffusion even at low current densities. Figure 8 (b1) and (b2) present the CO2 concentration profiles within the cell during discharge at i=1 mA/cm2 and i=0.1 mA/cm2, respectively. Figure S4 compares concentration profiles at various current densities. At i=1 mA/cm2, the CO2 concentration () decreases sharply from its initial value of 4.8 mol/m3 to 0 mol/m3 in the porous cathode, indicating strongly impeded CO2 diffusion. At i=0.1 mA/cm2, the decrease is more gradual. The impediment to CO2 transport, which is inversely proportional to the current density, results from the accumulation of Li2CO3 and C on the active cathode surface. This accumulation reduces pore availability for the electrolyte, increases the amount of insulating species, and obstructs pores, collectively restricting CO2 transport and limiting cell capacity (Figure 6 (c)). The nucleation and growth of Li2CO3 significantly impact performance, as described in Figure 4 and the detailed FEA model in section 2.4. Current density plays a crucial role in this process. At high current densities (1 mA/cm2), rapid nucleation near the CO2 feed side forms a thin Li2CO3/C film (Figure 8 (a1), (b1)). In contrast, low current densities (0.1 mA/cm2) promote slower growth, resulting in larger, more distributed Li2CO3 structures and a gradual decrease in porosity (Figure 8 (a2), (b2)).
To address these limitations, an ideal cathode should feature: a large surface area for reaction sites and product deposition, large pores to facilitate active species transport and prevent clogging, high porosity to accommodate discharge product growth, and high electrical conductivity and stability. Such a cathode design could mitigate the observed limitations and potentially enhance the overall performance of Li-CO2 batteries [77]. Finally, we note that direct experimental validation is essential to confirm our model’s mechanistic predictions. To rigorously assess accuracy and applicability, high-resolution imaging techniques such as transmission electron microscopy (TEM) and scanning electron microscopy (SEM) should be used to verify predicted changes in Li2CO3/C morphology, cathode surface features, and Li2CO3/C aggregation states. Additionally, time-resolved or ex-situ X-ray diffraction (XRD) can track the evolution of Li2CO3/C crystal structure and phase composition, while gas adsorption measurements, including Brunauer-Emmett-Teller (BET) and Barrett-Joyner-Halenda (BJH) analysis, can provide a quantitative assessment of cathode porosity and specific surface area. Importantly, a multiscale experimental approach is also needed to capture material behavior across different length and time scales, ensuring that both nanoscale mechanisms and macroscopic properties are accurately linked and described. These targeted and multiscale experiments are crucial for establishing the model’s predictive accuracy and real-world relevance.
4 Conclusion
Our interactive multiscale modeling framework, integrating FEA, MD, and DFT/AIMD, provides comprehensive insights into Li-CO2 battery electrochemistry. This approach enables accurate prediction of cell performance by considering critical factors such as CO2 transport, cathode porosity, and Li2CO3/C deposition.
DFT/AIMD calculations reveal the high electrical conductivity of the cathode catalyst, while MD simulations demonstrate favorable electrolyte properties. FEA results show an exponential decrease in discharge capacity with increasing current density-from 81,570 mAh/g at 0.1 mA/cm2 to 6,200 mAh/g at 1 mA/cm2-primarily due to CO2 transport limitations and reduced cathode porosity. The morphology of Li2CO3/C deposits transitions from larger structures at lower current densities to thin films at higher current densities, significantly impacting performance.
Our multiscale approach bridges atomic-level phenomena and cell-level performance, offering unique insights into the spatiotemporal evolution of key battery parameters and reducing the need for extensive experimental prototyping. Future research should focus on enhancing CO2 transport, improving discharge product decomposition, and optimizing cathode design with large surface areas, optimal pore sizes, and high porosity. This framework accelerates the development of high-performance Li-CO2 batteries and sustainable energy storage technologies by providing insights into phenomena that are difficult to observe experimentally in real time. Moreover, the developed Li-CO2 battery continuum model at the cell level can be extended to study battery packs and thermal runaway. This modeling framework can also be adapted to other battery chemistries or used to simulate degradation, cycling behavior, and thermal effects, thereby broadening its applicability.
CRediT Authorship Contribution Statement
Mohammed Lemaalem, Selva C. Selvaraj, and Naveen K. Dandu: Conceptualization, Visualization, Software, Data Curation, Formal Analysis, Methodology, Writing – Original Draft, Writing – Review & Editing. Ilias Papailias and Arash Namaeighasemi: Methodology, Data Curation, Formal Analysis, Writing – Review & Editing. Larry A. Curtiss and Amin Salehi-Khojin: Supervision, Data Curation, Formal Analysis, Writing – Review & Editing. Anh T. Ngo: Supervision, Conceptualization, Data Curation, Formal Analysis, Investigation, Project Administration, Writing – Review & Editing.
Author Contributions
This work was conceived by A. T. N.. M. L., S. C. S., and N. K. D., carried out the calculations, including first-principles DFT, AIMD results, large-scale MD simulations, and Finite Element Analysis. I. P., A. N., L. A. C., and A. S. K. were responsible for the experimental measurements. All authors contributed to the discussion of the results and the final manuscript.
Declaration of Competing Interest
The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.
Acknowledgements
This work was supported by the Assistant Secretary for Energy Efficiency and Renewable Energy, Office of Vehicle Technologies of the US Department of Energy, through the Battery Materials Research (BMR) program. We gratefully acknowledge the computing resources provided on Bebop, a high-performance computing cluster operated by the Laboratory Computing Resource Center (LCRC) at Argonne National Laboratory.
Appendix A. Supplementary material
Background on catalyst/electrolyte developments for Li-CO2 battery; Description of the experimental setup; Computational details; Plots showing the voltage-capacity curve during Li-CO2 battery discharge at various rates ranging from 0.2 mA/cm2 to 0.9 mA/cm2; Plots showing the effect of initial values of relevant variables on the voltage-capacity curve during Li-CO2 battery discharge at a rate of 0.4 mA/cm2; Plots illustrating changes in porosity and volume fraction of deposited species in the cathode throughout the discharge process at different current densities and discharge states; and Plots showing the change in CO2 concentration within the cathode.
Data availability
The data that support the plots within this paper and other findings of this study are available from the corresponding author upon reasonable request.